OpenStudy (anonymous):

Rationalize the denominator square root 3 over sq.root26+5 does anyone know how to figure this out?

OpenStudy (anonymous):

Check out: http://www.helpalgebra.com/articles/rationalizedenominator.htm

OpenStudy (amistre64):

you have to multiply the top and bottom by sqrt(26+5) if that is in your denominator. What happens is the squareroots cancel and your left with 26+5 on the bottom. Actually, yo need to write the problem better cause it is kinda hard to tell what it is :)

OpenStudy (anonymous):

If this is your equation:\[\sqrt{3}/\left( \sqrt{26} + 5 \right)\], then you need to multiply (numerator & denominator) by what's called the conjugate radical of the denominator: \[\sqrt{26}-5\]

OpenStudy (amistre64):

mrslange knows it :)